How do I find the solution set, given the replacement set:
5x + 2y = 5 [(2, -2), (2, -7), (5, -4), (3, -5)}

Answer by praseena(37)

5x + 2y = 5
The replacement set is [(2, -2), (2, -7), (5, -4), (3, -5)}
At first input the first set of values, that is (2,-2) in the given equation,
==> 5*2 + 2*-2 = 10 + -4 = 6
That is the value we got while substituting in the left side of the given equation is not same as the right side.
So (2,-2) is not a solution for the given equation.
===============================================================================
Now substitute the second pair of values in the given equation.
==> 5 * 2 + 2 * -7 = 10 + -14 = -4, which is again not equal to 5
so (2,-7) is not a solution
==========================================================================
Now take (5,-4)
==> 5 * 5 + 2 * -4 = 25 + -8 = 16, which is again not equal to 5
so this is also not a solution for the given equation.
=============================================================================
Now take the pair of values, (3,-5)
==> 5*3 + 2*-5 = 15 + -10 = 5, which is equal to the right side.
so (3,-5) is the solution set of the given equation.
O.K
BYE.